# Fusion Complexity Inversion: Why Simpler Cross View Modules Outperform SSMs and Cross View Attention Transformers for Pasture Biomass Regression

### 08 Mar 2026 v1

@@ III-C Architecture
Fusion: Gated Depthwise Convolution. The concatenated token sequence is processed by two stacked GatedDepthwiseConvBlocks. Each block applies LayerNorm, sigmoid gating, depthwise 1D convolution (), linear projection with dropout (), and a residual connection:
@@ IV-F The Metadata Paradox
| Fusion | No Meta | With Meta | |
@@ V Discussion
Limitations. All findings are derived from a single dataset, though no comparable alternative with laboratory validated, component wise biomass ground truth for proximal pasture imagery currently exists. The fusion complexity inversion may not hold where global modules have sufficient data to avoid overfitting. The backbone comparison partially confounds pretraining data size with algorithmic improvements, and 8 GB VRAM limits batch sizes to 1. Validation on emerging agricultural benchmarks with similar ground truth quality is left to future work.
@@ VI Conclusion
Foundation model adaptation for agricultural imagery is systematically studied on the 357 image CSIRO Pasture Biomass benchmark through 17 configurations. Three findings are established. First, fusion complexity inversion: a two layer gated depthwise convolution () outperforms cross view attention transformers (), bidirectional SSMs (), and full SSMs (, below the no fusion baseline). Second, foundation model dominance: rises monotonically from EfficientNet-B3 () to DINOv3-ViT-L (), confirming representation quality as the primary bottleneck. Third, the metadata fusion trap: training only metadata creates a ceiling at , collapsing an 8.4 point spread to 0.1 points. For scarce agricultural data, backbone quality should be prioritized, local fusion preferred, and inference-unavailable modalities excluded.

### 11 Apr 2026 v3

@@ III-C Architecture
Fusion: Gated Depthwise Convolution. The concatenated token sequence is processed by two stacked GatedDepthwiseConvBlocks, combining depthwise separable convolution [38] with multiplicative gating [37]. Each block applies LayerNorm, sigmoid gating, depthwise 1D convolution (), linear projection with dropout (), and a residual connection:
@@ IV-F The Metadata Paradox
| Fusion | No Meta () | With Meta () | |
@@ V Discussion
Limitations. All findings are derived from a single dataset, though no comparable alternative with laboratory validated, component wise biomass ground truth for proximal pasture imagery currently exists. The fusion complexity inversion may not hold with sufficient data, and the backbone comparison partially confounds pretraining data size with algorithmic improvements. At larger scales (like 2K–10K labeled images), complex global fusion modules would likely catch up to, or exceed local fusion. The 8 GB VRAM limit constrains effective batch size and training throughput. Validation on emerging agricultural benchmarks with similar ground truth quality is left to future work.
@@ VI Conclusion
Foundation model adaptation for agricultural imagery is systematically studied on the 357 image CSIRO Pasture Biomass benchmark through 17 configurations. Three findings are established. First, fusion complexity inversion: a two layer gated depthwise convolution () outperforms cross view attention transformers (), bidirectional SSMs (), and full SSMs (, below the no fusion baseline). Second, foundation model dominance: rises monotonically from EfficientNet-B3 () to DINOv3-ViT-L (), confirming representation quality as the primary bottleneck. Third, the metadata fusion trap: training only metadata creates a ceiling at , collapsing an 8.4 point spread to 0.1 points. For scarce agricultural data, backbone quality should be prioritized, local fusion preferred, and inference-unavailable modalities excluded. Code: https://github.com/WhiteMetagross/FusionComplexityInversionBiomass
Acknowledgment
We gratefully acknowledge the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for the creation, and public release of the CSIRO Pasture Biomass dataset, which made this research possible. Additionally, we disclose the use of AI assisted technologies: we utilized GPT 5.3 Codex, Claude Sonnet 4.6, and Claude Opus 4.6 for assistance with code implementation, grammar refinement, and writing and formatting support. Following the use of these tools, we thoroughly reviewed and edited the generated material as needed. We assume full responsibility for all scientific judgments, experimental design choices, and the final content of the manuscript.